OpenStudy (anonymous):

A punter kicks a football upward with an initial speed of 48 feet per second. After how many seconds does the ball hit the ground? Use the formula h = rt − 16t2 where h represents height in feet and r represents the initial speed.

OpenStudy (anonymous):

\[Vi=48m/s\]\[g=9.81m/s \]g=downward soo you h=\[48m/s \div9.81m/s=4.892...\]

OpenStudy (anonymous):

now you need to rearrange ten formula for t is that formula saying\[h=rt-16t^2\]?

OpenStudy (anonymous):

@LRGAUS95 ?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

@timo86m do u kno this

OpenStudy (anonymous):

h = rt − 16t2 0= rt − 16t2

OpenStudy (anonymous):

3 seconds ooh r is 48

OpenStudy (anonymous):

so the answer is 3 seconds

OpenStudy (anonymous):

@ConradBarczyk you dont have to do 9.8 m/s^2 the formula is done for you :)

OpenStudy (anonymous):

Yes 3 seconds

OpenStudy (anonymous):

LOL XD good point lol i did find the rest of the hard way :P \[t=\sqrt{h \div -16 +r}\]